rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#4074 — FS#8012 — Fiber between PLA93 and COU93

Attached to Project— ADSL / SDSL
Incident
Backend / Core
CLOSED
100%
Following a partial cleavage of cable this morning, which did not impact us, the teams are on the field had to cut the cable 144 FO in order to switch all on the new cable.

The traffic is not affected by this cut-off.

Expected delay for restoration: 3h
Date:  Wednesday, 06 February 2013, 07:11AM
Reason for closing:  Done
Comment by OVH - Wednesday, 06 February 2013, 07:10AM

Welds completed.
The loop is secured again.